ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 106, 1998

Amberg, B. and L. S. Kazarin On finite products of soluble groups  93

Cepedello Boiso, M. Approximation of Lipschitz functions by Δ-convex functions in Banach spaces

269

De León, M. See M. Fernández

Dix, J. G. Infinite frequency solutions to discontinuous control systems with variable delay

165

Fernández, M., R. Ibáñez and M. de León The canonical spectral sequences for Poisson manifolds

133

Hájek, P. Polynomial algebras on classical Banach spaces  209

Ibáñez, R. See M. Fernández

Jiménez-Seral, P. and J. P. Lafuente On complemented nonabelian chief factors of a finite group

177

Jonoska, N. A conjugacy invariant for reducible sofic shifts and its semigroup characterizations

221

Kamae, T. Linear expansions, strictly ergodic homogeneous cocycles and fractals

313

Kazarin, L. S. See B. Amberg

Koldobsky, A. An application of the Fourier transform to sections of star bodies

157

Lafuente, J. P. See P. Jiménez-Seral

Ledet, A. Embedding problems with cyclic kernel of order 4  109
Leung, K. H. and S. H. Man On Schur rings over cyclic groups  251

Linial, N., A. Magen and M. E. Saks Low distortion Euclidean embeddings of trees

339

Magaard, K., K. Strambach and H. Völklein Finite quotients of the pure symplectic braid group

13

Magen, A. See N. Linial

Man, S. H. See K. H. Leung

Morris, G. and T. Ward Entropy bounds for endomorphisms commuting with K actions

1

Newelski, L. M-gap conjecture and m-normal theories  285

Saks, M. E. See N. Linial Strambach, K. See K. Magaard Völklein, H. See K. Magaard Ward, T. See G. Morris

Wasserman, A. A derivation HNN construction for Lie algebras  79

Weiss, B. Finite dimensional representations and subgroup actions on homogeneous spaces

189

Zinoviev, D. Relation of orbital integrals on SO(5) and PGL(2)  29